# Electric field tunable magnetoexcitons in Xenes-hBN-TMDC, Xenes-hBN-BP, and Xenes-hBN-TMTC heterostructures

### v1

@@ II Theoretical Model
Figure 2: (Color online) The dependence of the reduced mass of an exciton in Xenes/TMDC ($a$) and Xenes/phosphorene ($b$) heterostructures on the electric field.
@@
III Results of Calculations and Discussion
First, let us focus on the reduced mass in the Xenes-hBN-TMDC and Xenes-hBN-BP heterostructures. The dependence of the exciton reduced mass in these heterostructures is depicted in Fig. 2. Analysis of the results in Fig. 2 reveals a distinct behavior of the reduced masses. In the Xenes-hBN-TMDC heterostructure, the exciton reduced mass is smaller than that in the WS2 for electric fields $E_{\perp}<1.5$ V/m, but becomes larger for $E{\perp}>1.5$ V/m. In contrast, the reduced masses $\mu_{x}$ and $\mu_{y}$ in the Xenes-hBN-BP heterostructure remain smaller than those in phosphorene across the entire range of electric field values.
@@ III Results of Calculations and Discussion
The relationship between BE and reduced mass has significant implications in monolayer physics. In TMDC-hBN-TMDC heterostructures, which exhibit lower electron and hole effective masses than Xenes-hBN-TMDC at $E_{\perp}>1.5$ V/m, excitons tend to have lower binding energies. This characteristic behavior affects the efficiency of exciton-based devices, as lower binding energies facilitate the dissociation of excitons into free charge carriers—an essential process for efficient charge transport. Understanding this dependence is crucial for the design of materials with tailored electronic and optical properties, particularly in systems where excitonic effects play a dominant role.
Figure 4: (Color online) The dependence of the total potentials $W_{1}(x,y)=\frac{e^{2}}{8\mu_{x}}B^{2}x^{2}+\frac{e^{2}}{8\mu_{y}}B^{2}y^{2}+V(x,y)$ and $W_{2}(x,y)=\frac{e^{2}}{8\mu}B^{2}(x^{2}+y^{2})+V(x,y)$ acting on the electron-hole pair in Xenes-hBN-phosphorene and Xenes-hBN-TMDC ($a$) and Xenes-hBN-TMTC and Xenes-hBN-TMDC ($b$) heterostructures. Calculations are performed for the electron-hole pair in the external magnetic field $B=30$ T and electric field $E_{\perp}=0.5$ V/m.
Figure 5: (Color online) Dependence of energy $\Delta E$ for magnetoexcitons in $1s$ and $3s$ states in Xenes/MoSe2 and Xenes-hBN-BP heterostructures on the electric and magnetic fields.
In the next step, let us consider indirect magnetoexcitons formed by electron-hole pair in van der Waals heterostructures (e.g., Xenes–hBN–TMDC, Xenes–hBN–BP, Xenes–hBN–TMTC), in the presence of magnetic field. In these vdW heterostructures magnetoexcitons emerge as a result of the influence of an external magnetic field on excitons, i.e. the presence of an external magnetic field alters the excitonic states and leads to the formation of magnetoexcitons, which are excitons whose properties are modified by the magnetic field. The characteristics of magnetoexcitons are governed by both the underlying electronic structure of the heterostructure and the strength of the magnetic field, modifying their internal structure and energy spectrum. In Eqs. (5) and (6) $W_{1}(x,y)=\frac{e^{2}}{8\mu_{x}}B^{2}x^{2}+\frac{e^{2}}{8\mu_{y}}B^{2}y^{2}+V(x,y)$ and $W_{2}(x,y)=\frac{e^{2}}{8\mu}B^{2}(x^{2}+y^{2})+V(x,y)$ are total potentials that act on the electron-hole pair in these heterostructures, respectively. The Coulomb interaction $V(x,y)$ (3) as well as $W_{2}(x,y)$ have a rotational symmetry. In contrast, though, the terms $\frac{e^{2}}{8\mu_{x}}B^{2}x^{2}$ and $\frac{e^{2}}{8\mu_{y}}B^{2}y^{2}$ in $W_{1}(x,y)$ break the rotational symmetry due to its anisotropic nature. As a result of reduced masses anisotropy, when $\mu_{x}>\mu_{y}$, $\frac{e^{2}}{8\mu_{y}}B^{2}y^{2}$ gives larger contribution due to the magnetic field than the term $\frac{e^{2}}{8\mu_{x}}B^{2}x^{2}$ and vice versa if $\mu_{y}>\mu_{x}$. Thus, the electron and hole masses anisotropy leads to the anisotropy of $W_{1}(x,y)$. The dependence of the total potential $W_{1}(x,y)$ and $W_{2}(x,y)$, respectively, acting on the electron-hole pair on $x$ and $y$ coordinates is shown in Fig. 4. Let us emphasize that the anisotropy of the potential $W_{1}(x,y)$ with respect to $W_{2}(x,y)$ is different in the case of heterostructures with phosphorene and TMTC.
The results of calculations of the energy contribution $\Delta E$ due to a magnetic field at different values of the electric field are presented in Fig. 5. Here, we consider MoSe2 and phosphorene as representative cases for isotropic and anisotropic materials, respectively. Contributions of the magnetic field are proportional to the factors $1/\mu$ for an isotropic material and $1/\mu_{x}$ and $1/\mu_{y}$ for an anisotropic material. Therefore, the electric field increase leads to the decrease of $1/\mu$, $1/\mu_{x}$, and $1/\mu_{y}$, consequently, $\Delta E$ decreases. At the small electric field $1/\mu$, $1/\mu_{x}$, and $1/\mu_{y}$ are comparable, although $1/\mu_{x}$, and $1/\mu_{y}$ a little bit exceed $1/\mu$. By increasing the number of hBN layers, the electrostatic attraction decreases, which changes $\Delta E$. When $N=1$ the difference between $\Delta E$ for both heterostructures is negligible (Fig. 5 $a$ $b$). When $E_{\perp}>1$ V/m, $1/\mu<1/\mu_{x}$ and $1/\mu<1/\mu_{y}$, and $\Delta E$ becomes larger for heterostructure with phosphorene (Fig. 5 $c$ $f$). In Fig. 5, we demonstrate this for $1s$ and $3s$ states, but the same tendencies can be extended to $2s$ and $4s$ states. In Xenes–hBN–MoSe2 and Xenes–hBN–BP heterostructures, the strong field-tunable band gaps and anisotropic effective masses of phosphorene can lead to pronounced shifts in exciton binding energy and modified magnetoexciton behavior compared to conventional heterostructures build with TMDC or phosphorene monolayers. Due to the larger effective dielectric constant introduced by hBN layers, excitons are more weakly bound, making their magneto-optical response more sensitive to external fields.
The diamagnetic coefficients for excitons in Si-hBN-MoSe2 and Si-hBN-BP heterostructures, obtained in the framework of our approach, are reported in Table 1. For Xenes-hBN-TMDC we show a representative case - Si-hBN-MoSe2. Analysis of the results presented shows: i. the diamagnetic coefficients are decreasing with increasing electric field for all Rydberg states in both heterostructures; ii. the increase in the number of insulating hBN layers leads to the increase of $\sigma$ for $1s$, $2s$ and $3s$ states; iii. $3s$ and $4s$ states generally exhibit high diamagnetic coefficients due to the increase exciton’s $r^{2}$ affected by the size of the exciton wavefunction.
Table 1: Diamagnetic coefficients for magnetoexcitons in Si-hBN-phosphorene and Si-hBN-MoSe2. DMCs are obtained when $R^{2}=0.9998$ for the linear regression model.
Diamagnetic coefficient $\gamma_{2}$ [$\mu$eV/T2]
@@
$H_{\bf p}(t)=\begin{pmatrix}p^{2}/2\mu(t)&\Delta_{p}\\ \Delta_{p}&-p^{2}/2\mu(t)\end{pmatrix},$ (10)
where $\Delta_{p}$ represents the uniform band splitting due to electron-hole pairing and $\mu(t)$ the time-dependent reduced mass. $\Delta_{p}$ can be calculated self-consistently, following the approach of Ref. [22], where the mass oscillations are corrections in comparison to the time-average mass. At low density, $\Delta_{p}>0$ for a sufficiently strong interlayer attraction between electrons and holes. The other condition is that the gap $2\Delta_{p\xi\sigma}=2|\xi\sigma\Delta_{so}-ed_{0}E_{0}\cos wt|$ must be larger than the kinetic energy $p^{2}/2\mu$. This means that $E_{0}$ must be smaller than $\Delta_{so}/ed_{0}$. Thus, we can assume within the leading approximation that only the reduced mass $\mu(t)$ depends on the external electric field. From $H_{\bf p}(t)$ we obtain the Floquet Hamiltonian [62] as
@@
Its eigenvalue spectrum yields the effective band structure of the Floquet approach. This can be understood in the present case as if the variation of the reduced mass $\mu(t)$ increases and decreases the bandwidth periodically in time. Since the Floquet Hamiltonian is created by the unitary operator, values of the spectrum of $H_{\bf p}(t)$ that are outside the interval $[-\pi,\pi]$ are backfolded on the complex unit circle. This leads to the restriction of the bandwidth to $2\pi$. This effect is illustrated in Fig. 6, which reveals that the time-dependent electric field applied to the layered system provides a tool for band-structure engineering, similar to the concept presented in Ref. [63].
Figure 6: (Color online) Typical effect of Floquet band engineering: while ($a$) represents the static case, ($b$) is the result of the Floquet dynamics due to a periodically time-dependent exciton mass, which represents a backfolded spectrum on the interval $[-\pi,\pi]$.
@@ IV Concluding remarks
In this work we suggest the novel vdW heterostructures, comprising Xenes, TMDCs, phosphorene, and TMTCs monolayers separated by hBN insulating layers. We have investigated theoretically the properties of Rydberg indirect excitons in Xenes-hBN-TMDC, Xenes-hBN-BP, and Xenes-hBN-TMTC heterostructures subjected to parallel external electric and magnetic fields that are oriented perpendicular to the layers. By incorporating both isotropic and anisotropic materials, we demonstrated the ability to control and tune excitonic properties through external fields and structural parameters.
@@
In summary, vdW heterostructures, composed of two layers of 2D materials when one of the layers is Xenes, offer a versatile platform for exploring novel electronic and optical properties. By carefully engineering the stacking, orientation, and strain of these materials and applying external electric and magnetic fields, researchers can design systems with tailored properties, which are suitable for a wide range of applications in next-generation electronic and optoelectronic devices.

### v2

@@ II Theoretical Model
Figure 2: (Color online) The dependence of the reduced mass of an exciton in Xenes/TMDC ($a$) and Xenes/BP ($b$) heterostructures on the electric field. In calculations, we used masses of holes in WS2 and BP monolayers from [67] and [69], respectively.
Figure 3: (Color online) The electric-field dependence of the binding energies of an exciton in the Rydberg state 1ss for $(a)$ Si/hBN/MoSe2 and $(b)$ Si/hBN/BP heterostructures. The number of dielectric layers, hBN, is given for NN = 1, 6. In calculations, we used masses of holes and thicknesses of MoS2 and BP monolayers from [64] and [67], and [69] and [72], respectively.
III Results of Calculations and Discussion
First, let us focus on the reduced mass in the Xenes/hBN/TMDC and Xenes/hBN/BP heterostructures. There is interesting physics here, but the presentation can benefit from clarification. The magnetic contribution to the exciton binding energy is inversely proportional to the exciton’s reduced mass, which is always smaller than the smaller of the two constituent masses. In TMDCs and phosphorene, the hole effective mass typically exceeds that of the electron. In contrast, Xenes exhibit equal and electrically tunable electron and hole effective masses ($m^{e}=m^{h}$). Consequently, by localizing the hole in BP, TMDC, or TMTC layer, one can achieve a substantial variation in the reduced mass—potentially reaching values smaller than the electron effective mass in TMDCs or BP.
This dependence of the reduced mass on the heterostructure composition is illustrated in Fig. 2. The analysis reveals that the reduced mass in Xenes/TMDC and Xenes/BP heterostructures is smaller than in TMDC/TMDC or BP/BP counterparts. As follows from Fig. 2, this reduction of mass enhances the magnetic contribution to the binding energy and increases its tunability with an applied electric field.
For the calculations of reduced masses, we used Eqs. (9) and (7) for isotropic and anisotropic heterostructures, respectively. The effective masses of electrons and holes in TMDC materials were obtained using various methods [63]. In Xenes/hBN/TMDC heterostructures, we used hole effective mass values from Refs. [64, 65, 66, 63], while the electron mass was calculated using Eq. (3).
In calculations for the Xenes/hBN/BP heterostructure, the effective masses of electrons and holes can be taken from Refs. [68, 69, 70, 71, 72], which are based on first-principles calculations. However, the lattice constants reported in these references differ, as do the exchange-correlation functionals and tight-binding parameter sets used in the simulations. These variations naturally lead to discrepancies in the anisotropic effective masses, particularly in the curvatures of the conduction and valence bands along the armchair and zigzag directions. In calculations for Xenes/hBN/TMTC heterostructures, one can utilize the data for effective electrons and holes masses reported in [20, 73, 74, 75] that are summarized in [8]. In our calculations, we use input parameters listed in Table 1.
Ab initio calculations [37] indicate that the crystal structure of silicene becomes unstable at electric fields around 2.6 V/Å. The band gap in Xenes can be tuned by an external electric field up to this critical strength. However, in the proposed van der Waals heterostructures, the Xene layer is separated from the adjacent 2D monolayer by insulating hBN spacers. The electric breakdown field of hBN monolayer is approximately 0.1 V/Å, although it depends on both the thickness and the number of hBN layers[77, 78]. A recent study [79] reports that, under certain conditions, the breakdown threshold for a monolayer can be increased to about 0.4 V/Å. Accordingly, in our calculations, we considered electric field strengths in the range of 0.1-0.5 V/Å.
Table 1: Input parameters. Reduced masses in units of the mass of electron $m_{0}$
Xenes
@@ III Results of Calculations and Discussion
Si 38 [76] 0.46 [76] 5.06 [76] 3.33 [76]
BP and TMTC
$\mu_{x}^{e}$ $m_{0}$] $\mu_{y}^{h}$ $m_{0}$] $h$ [Å]
BP 0.64496 [69] 0.20488 [69] 5.41 [72]
ZrS3 1.26 [20] 0.33 [20] 8.9 [20]
TMDC
$\mu$ $m_{0}$] $h$ [Å]
WS2 0.23 [64] 6.219 [67]
MoSe2 0.31 [64] 6.527 [67]
We numerically solve Eqs. (6) and (8) using a code implemented in Refs. [8, 31, 33, 34, 35]. The method utilizes the finite element method implemented in Wolfram Mathematica in the NDEigensystem function, which yields pairs of eigenenergies and eigenfunctions corresponding to the most strongly bound states. Parameters are taken from above above-mentioned references and summarized in Table 1. Notice that parameters are not as important since in the literature for each material, there is a range for each parameter. Therefore, our work demonstrates the viability of the systems rather than exact numbers. The results change based on the parameters used.
As an illustrative example, we present below the results for $A$-type indirect excitons. Including $B$-type excitons does not lead to any qualitative changes in the conclusions.
The relationship between the binding energy (BE) and the reduced mass of an exciton arises from the nature of the Coulomb interaction between the electron and the hole. The exciton binding energy, defined as the energy required to dissociate the exciton into free electron and hole, is proportional to the reduced mass of the electron–hole system. Thus, a larger reduced mass leads to a stronger binding energy. Consequently, as the electric field increases, the reduced mass also increases, resulting in a corresponding increase in the BE. This dependence is illustrated in Fig. 3 for excitons in the Rydberg state 1ss. Also, from Fig. 3, one can conclude that increasing the number of hBN layers results in a notable reduction in the exciton binding energy due to enhanced dielectric screening effects.
The relationship between BE and reduced mass has significant implications in monolayer physics. In TMDC/hBN/TMDC heterostructures, which exhibit higher reduced masses than Xenes/hBN/TMDC, excitons tend to have higher binding energies. This characteristic behavior affects the efficiency of exciton-based devices, as lower binding energies facilitate the dissociation of excitons into free charge carriers—an essential process for efficient charge transport. Understanding this dependence is crucial for the design of materials with tailored electronic and optical properties, particularly in systems where excitonic effects play a dominant role.
Figure 4: (Color online) The dependence of the total potentials $W_{1}(x,y)=\frac{e^{2}}{8\mu_{x}}B^{2}x^{2}+\frac{e^{2}}{8\mu_{y}}B^{2}y^{2}+V(x,y)$ and $W_{2}(x,y)=\frac{e^{2}}{8\mu}B^{2}(x^{2}+y^{2})+V(x,y)$ acting on the electron-hole pair in Xenes/BP (Si/BP) and Xenes/TMDC (Si/WS2) ($a$) and Xenes/TMTC (Si/ZrS3) and Xenes/TMDC (Si/WS2) $b$) heterostructures. Calculations are performed for the electron-hole pair in the external magnetic field $B=30$ T and electric field $E_{\perp}=0.1$ V/Å.
Figure 5: (Color online) Dependence of energy $\Delta E$ for magnetoexcitons in $1s$ and $3s$ states in Xenes/MoSe2 and Xenes/hBN/BP heterostructures on the electric and magnetic fields. In calculations, we used parameters summarized in Table 1.
In the next step, let us consider indirect magnetoexcitons formed by electron-hole pair in van der Waals heterostructures (e.g., Xenes/hBN/TMDC, Xenes/hBN/BP, Xenes/hBN/TMTC), in the presence of a magnetic field. In these vdW heterostructures magnetoexcitons emerge as a result of the influence of an external magnetic field on excitons, i.e. the presence of an external magnetic field alters the excitonic states and leads to the formation of magnetoexcitons, which are excitons whose properties are modified by the magnetic field. The characteristics of magnetoexcitons are governed by both the underlying electronic structure of the heterostructure and the strength of the magnetic field, modifying their internal structure and energy spectrum. In Eqs. (6) and (8) $W_{1}(x,y)=\frac{e^{2}}{8\mu_{x}}B^{2}x^{2}+\frac{e^{2}}{8\mu_{y}}B^{2}y^{2}+V(x,y)$ and $W_{2}(x,y)=\frac{e^{2}}{8\mu}B^{2}(x^{2}+y^{2})+V(x,y)$ are total potentials that act on the electron-hole pair in these heterostructures, respectively. The Coulomb interaction $V(x,y)$ (4) as well as $W_{2}(x,y)$ have a rotational symmetry. In contrast, though, the terms $\frac{e^{2}}{8\mu_{x}}B^{2}x^{2}$ and $\frac{e^{2}}{8\mu_{y}}B^{2}y^{2}$ in $W_{1}(x,y)$ break the rotational symmetry due to its anisotropic nature. As a result of reduced masses anisotropy, when $\mu_{x}>\mu_{y}$, $\frac{e^{2}}{8\mu_{y}}B^{2}y^{2}$ gives larger contribution due to the magnetic field than the term $\frac{e^{2}}{8\mu_{x}}B^{2}x^{2}$ and vice versa if $\mu_{y}>\mu_{x}$. Thus, the electron and hole masses anisotropy leads to the anisotropy of $W_{1}(x,y)$. The dependence of the total potential $W_{1}(x,y)$ and $W_{2}(x,y)$, respectively, acting on the electron-hole pair on $x$ and $y$ coordinates is shown in Fig. 4. Let us emphasize that the anisotropy of the potential $W_{1}(x,y)$ with respect to $W_{2}(x,y)$ is different in the case of heterostructures with phosphorene and TMTC.
The results of calculations of the energy contribution $\Delta E$ due to a magnetic field at different values of the electric field are presented in Fig. 5. Here, we consider MoSe2 and phosphorene as representative cases for isotropic and anisotropic materials, respectively. Contributions of the magnetic field are proportional to the factors $1/\mu$ for an isotropic material and $1/\mu_{x}$ and $1/\mu_{y}$ for an anisotropic material. Therefore, the electric field increase leads to the decrease of $1/\mu$, $1/\mu_{x}$, and $1/\mu_{y}$, consequently, $\Delta E$ decreases. At the small electric field $1/\mu$, $1/\mu_{x}$, and $1/\mu_{y}$ are comparable, although $1/\mu_{x}$, and $1/\mu_{y}$ a little bit exceed $1/\mu$. By increasing the number of hBN layers, the electrostatic attraction decreases, which changes $\Delta E$. In Fig. 5, we demonstrate this for $1s$ and $3s$ states, but the same tendencies can be extended to $2s$ and $4s$ states. In Xenes/hBN/MoSe2 and Xenes/hBN/BP heterostructures, the strong field-tunable band gaps and anisotropic effective masses of phosphorene can lead to pronounced shifts in exciton binding energy and modified magnetoexciton behavior compared to conventional heterostructures built with TMDC or phosphorene monolayers. Due to the larger effective dielectric constant introduced by hBN layers, excitons are more weakly bound, making their magneto-optical response more sensitive to external fields.
The diamagnetic coefficients for excitons in Si/hBN/MoSe2 and Si/hBN/BP heterostructures, obtained in the framework of our approach, are reported in Table 2. For Xenes/hBN/TMDC we show a representative case - Si/hBN/MoSe2. Analysis of the presented results shows: i. the diamagnetic coefficients are decreasing with increasing electric field for all Rydberg states in both heterostructures; ii. the increase in the number of insulating hBN layers leads to the increase of $\gamma_{2}$ for $1s$, $2s$ and $3s$ states; iii. $3s$ and $4s$ states generally exhibit high diamagnetic coefficients due to the increase exciton’s $r^{2}$ affected by the size of the exciton wavefunction.
Table 2: Diamagnetic coefficients for magnetoexcitons in Si/hBN/BP and Si/hBN/MoSe2. DMCs are obtained when $R^{2}=0.9998$ for the linear regression model. Some cells are empty because diamagnetic coefficients could not be extracted with linear regression satisfying $R^{2}=0.9998$.
@@
$H_{\bf p}(t)=\begin{pmatrix}p^{2}/2\mu(t)&\Delta_{p}(t)\\ \Delta_{p}(t)&-p^{2}/2\mu(t)\end{pmatrix}$, (13)
where $\Delta_{p}(t)$ represents the band splitting due to electron-hole pairing and $\mu(t)$ the time-dependent reduced mass. $\Delta_{p}(t)$ can be calculated self-consistently, following the approach of Ref. [22], where the mass oscillations are corrections in comparison to the time-average mass. At low density, $\Delta_{p}>0$ for a sufficiently strong interlayer attraction between electrons and holes. The other condition is that the gap $2\Delta_{p\xi\sigma}=2|\xi\sigma\Delta_{so}-ed_{0}E_{0}\cos wt|$ must be larger than the kinetic energy $p^{2}/2\mu$. This means that $E_{0}$ must be smaller than $\Delta_{so}/ed_{0}$. Thus, we can assume within the leading approximation that only the reduced mass $\mu(t)$ depends on the external electric field. From $H_{\bf p}(t)$ we obtain the Floquet Hamiltonian [80] as
@@
Its eigenvalue spectrum yields the effective band structure of the Floquet approach. This can be understood in the present case as if the variation of the reduced mass $\mu(t)$ increases and decreases the bandwidth periodically in time. Since the Floquet Hamiltonian is created by the unitary operator, values of the spectrum of $H_{\bf p}(t)$ that are outside the interval $\pi,\pi]$ are backfolded on the complex unit circle. This leads to the restriction of the bandwidth to $2\pi$. This effect is illustrated in Fig. 6, which reveals that the time-dependent electric field applied to the layered system provides a tool for band-structure engineering, similar to the concept presented in Ref. [81].
Figure 6: (Color online) Typical effect of Floquet band engineering: while ($a$) represents the static case, ($b$) is the result of the Floquet dynamics due to a periodically time-dependent exciton mass, which represents a backfolded spectrum on the interval $\pi,\pi]$. Since the Hamiltonian (13) is isotropic, the direction of the wavevector does not affect the spectrum.
@@ IV Concluding remarks
In this work we suggest the novel vdW heterostructures, comprising Xenes, TMDCs, phosphorene, and TMTCs monolayers separated by hBN insulating layers. We have investigated theoretically the properties of Rydberg indirect excitons in Xenes/hBN/TMDC, Xenes/hBN/BP, and Xenes/hBN/TMTC heterostructures subjected to parallel external electric and magnetic fields that are oriented perpendicular to the layers. By incorporating both isotropic and anisotropic materials, we demonstrated the ability to control and tune excitonic properties through external fields and structural parameters.
@@
In summary, vdW heterostructures, composed of two layers of 2D materials when one of the layers is Xenes, offer a versatile platform for exploring novel electronic and optical properties. By carefully engineering the stacking, orientation, and strain of these materials and applying external electric and magnetic fields, researchers can design systems with tailored properties.
Appendix A Approximation to decouple the relative and center-of-mass motions
The Schrödinger equation of an interacting eehh pair in a magnetic field $\mathbf{B}$ reads
$\left[\frac{1}{2m^{e}}\left(\mathbf{p}_{e}-e\mathbf{A}_{e}\right)^{2}+\frac{1}{2m^{h}}\left(\mathbf{p}_{h}+e\mathbf{A}_{h}\right)^{2}+V(|\mathbf{r}_{e}-\mathbf{r}_{h}|)\right]\Psi(\mathbf{r}_{e},\mathbf{r}_{h})=\mathcal{E}\,\Psi(\mathbf{r}_{e},\mathbf{r}_{h}),$ (A1)
where $\mathbf{p}_{e}=-i\hbar\nabla_{e}$ and $\mathbf{p}_{h}=-i\hbar\nabla_{h}$, $\mathbf{A}_{i}=-\mathbf{r}_{i}\times\mathbf{B}/2$ is the vector potential generating the magnetic field $\mathbf{B}$, and $V(|\mathbf{r}_{e}-\mathbf{r}_{h}|)$ is an electrostatic interaction. We assume the symmetric gauge corresponding to a uniform perpendicular external magnetic field $\mathbf{B}=(0,0,B)$.
Our focus is on Rydberg $s$-state excitons. The terms $e\mathbf{p}_{e}\mathbf{A}_{e}$ and $e\mathbf{p}_{h}\mathbf{A}_{h}$ in Eq. (A1) lead to an effective electric field that appears in the center-of-mass frame of the eehh system which interacts with the relative coordinate electric dipole $er$, and orbital momentum of the ehe-h. However, orbital momentum of the ehe-h pair for Rydberg excitons is zero. Since these terms are proportional to the magnetic field $B$, restricting ourselves to contributions proportional to $B^{2}$, we obtain
$\left[\frac{1}{2m^{e}}\left(\mathbf{p}^{2}_{e}+e^{2}\mathbf{A}_{e}^{2}\right)+\frac{1}{2m^{h}}\left(\mathbf{p}^{2}_{h}+e^{2}\mathbf{A}_{h}^{2}\right)+V_{ij}(|\mathbf{r}_{i}-\mathbf{r}_{j}|)\right]\Psi(\mathbf{r}_{e},\mathbf{r}_{h})=\mathcal{E}\,\Psi(\mathbf{r}_{e},\mathbf{r}_{h}).$ (A2)
At the next step we introduce the coordinates of the center-of-mass and relative motion for two particles. Making the assumption that the electron and hole have nearly equal masses, we obtain
$\bigg[-\frac{1}{2M}\frac{\partial^{2}}{\partial\mathbf{R}^{2}}-\frac{1}{2\mu}\frac{\partial^{2}}{\partial\mathbf{r}^{2}}+\frac{e^{2}}{8\mu}(\mathbf{B}\times\mathbf{R})^{2}+\frac{e^{2}}{8\mu}(\mathbf{B}\times\mathbf{r})^{2}+V(\mathbf{r})\bigg]\psi(\mathbf{R},\mathbf{r})=\mathcal{E}\,\psi(\mathbf{R},\mathbf{r}),$ (A3)
where $M=m^{e}+m^{h}$ is the total mass and $\mathbf{R}=(X,Y)$.
Finally, seeking the solution of Eq. (A3) in the form $\psi(\mathbf{R},\mathbf{r})=\Psi(\mathbf{R})\Phi(\mathbf{r})$, and separating the angular variable, the problem decouples into two equations describing the relative and the center-of-mass motion:
$\left[-\frac{1}{2\mu}\frac{\partial^{2}}{\partial r^{2}}-\frac{1}{2\mu}\frac{1}{r}\frac{\partial}{\partial r}+\frac{e^{2}}{8\mu}(\mathbf{B}\times\mathbf{r})^{2}+V(r)\right]\Phi(r)=E\Phi(r),$ (A4)
$\left[-\frac{1}{2M}\frac{\partial^{2}}{\partial\mathbf{R}^{2}}+\frac{e^{2}}{8\mu}(\mathbf{B}\times\mathbf{R})^{2}\right]\Psi_{cm}(\mathbf{R})=E_{cm}\Psi_{cm}(\mathbf{R}),$ (A5)
where $E$ and $E_{cm}$ are the eigenenergies of the relative and center-of-mass motions, respectively, so that $\mathcal{E}=E+E_{cm}$.
The Schrödinger equation for the center-of-mass (A5) formally corresponds to the isotropic harmonic oscillator with the effective mass $2\sqrt{M\mu}$ in 2D space, oscillated with frequency $\varpi=\frac{eB}{2\sqrt{M\mu}}$. In Cartesian coordinates it can be solved exactly, with the energy spectrum
$E_{cm}=\hbar\varpi\left(\mathfrak{n}_{X}+\mathfrak{n}_{Y}+1\right)=\hbar\varpi\left(\mathfrak{N}+1\right),$ (A6)
where $\mathfrak{N}=\mathfrak{n}_{X}+\mathfrak{n}_{Y}$, with $\mathfrak{n}_{X}=0,1,2,3,\ldots$ and $\mathfrak{n}_{Y}=0,1,2,3,\ldots$.
Thus, within the above-mentioned approximations, we can decouple the relative and the center-of-mass motions. The energy (A6) will contribute to the total energy of the magnetoexciton.